# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ A.1.2 Proof of Lemma 8
Proof of Lemma 10  Adamczak, (2015, Remark 2.2) shows that multinomial random vectors satisfy the convex concentration inequality (Adamczak,, 2015, Definition 2.2). Therefore, we can apply Adamczak, (2015, Theorem 2.5) to conclude that there exists constants $C_{1},C_{2}$ such that, for any $n\times n$ matrix $A$ and any $t>0$,
@@
[…]
@@
We are now ready to prove Lemma 8.
Proof of Lemma 8  Denote by $F_{\infty}$ the cumulative distribution function for the conditional distribution of the bootstrap sample $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$ (defined in (3)) given $\mathbb{X}_{n,\mathbf{z}}$, and denote its $(1-\alpha)$-quantile as
[…]
To bound $q^{2}_{\infty,1-\alpha/2}(\mathbb{X}_{n,\mathbf{z}})$, we use a similar argument as in Schrab et al., (2023, Appendix E4) by bounding the exceedance probability of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$. Notably, their proof cannot be applied directly here, because they use a wild bootstrap (Leucht and Neumann,, 2013) that corresponds to replacing $\mathbf{W}_{i}-1$ with independent Rademacher weights, whereas in our case $W_{i}$ are correlated and multinomially distributed. To proceed, we first note that we can write $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$ as a quadratic form as $D_{\mathbf{W}}^{2}(\mathbb{X}_{n,\mathbf{z}})=n^{-2}\mathbf{W}^{\top}A\mathbf{W}$, where $A=(a_{ij})_{ij}$ with $a_{ij}\coloneqq u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$. We can then use the Hanson-Wright inequality stated in Lemma 10 applied with $\delta=\alpha/2$ to show that there exist positive constant $C$ such that, for any almost sure realizations $\mathbb{X}_{n,\mathbf{z}}$,
@@
where the last quality holds since by assumption $\mathbf{X}_{i}=\mathbf{z}$ for $i>m$. Moreover, since $a_{ii}=u_{p}(\mathbf{X}_{i},\mathbf{X}_{i})\geq 0$ for any $i$, we can apply Markov inequality to bound the first term on the RHS. This gives that the following holds with probability at least $41-\delta/$ 4,
@@
where the second step holds since by assumption $\mathbf{X}_{i}=\mathbf{X}_{i}^{\ast}\sim P$ are i.i.d. for $i=1,\ldots,m$, and where we have substituted $\xi_{P}^{2}=\mathbb{E}_{\mathbf{X}_{1}\sim P}[u_{p}(\mathbf{X}_{1},\mathbf{X}_{1})^{2}]$. Combining with (15) implies that, with probability at least $41-\delta/$ 4,
@@
$\Sigma_{P}^{2}\;=\;\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]\;\leq\;\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X})u_{p}(\mathbf{X}^{\prime},\mathbf{X}^{\prime})]\;$ =$\stackrel{{\scriptstyle(a)}}{{=}}\;\big{(}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{X})]\big{)}^{2}$
@@
where $(a)$ holds since $\mathbf{X},\mathbf{X}^{\prime}$ are i.i.d. and $(b)$ follows from Jensen’s inequality. By applying Markov’s inequality again, with probability at least $41-\delta/$ 4, where the probability is taken over the randomness of $\mathbb{X}_{m}^{\ast}$, we have
@@
$q_{\infty,1-\alpha}^{2}(\mathbb{X}_{n,\mathbf{z}})\;$ $\leq\;\frac{4}{n\delta}\xi_{P}+\frac{n-m}{n^{2}}u_{p}(\mathbf{z},\mathbf{z})$
@@
A.1.3 Proof of Proposition 9
We first show that the moment conditions (i)-(ii) in Lemma 7 and Lemma 8 are met, and apply these lemmas to conclude the proof.
@@ A.1.3 Proof of Proposition 9
$u_{p}(\mathbf{x},\mathbf{z})^{2}\;$ $=\;\big{[}\mathbf{s}_{p}(\mathbf{x})^{\top}\mathbf{s}_{p}(\mathbf{z})h(\mathbf{x}-\mathbf{z})-\mathbf{s}_{p}(\mathbf{x})^{\top}\nabla h(\mathbf{x}-\mathbf{z})+\mathbf{s}_{p}(\mathbf{z})^{\top}\nabla h(\mathbf{x}-\mathbf{z})+\nabla^{\top}\nabla h(\mathbf{x}-\mathbf{z})\big{]}^{2}$
@@
$\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}\big{[}\|\mathbf{s}_{p}(\mathbf{X})\|_{2}^{2}\|\mathbf{s}_{p}(\mathbf{z})\|_{2}^{2}h(\mathbf{X}-\mathbf{z})^{2}\big{]}\;$ $\leq\;\sup_{\mathbf{z}\in\mathbb{R}^{d}}\|\mathbf{s}_{p}(\mathbf{z})\|_{2}^{2}\big{(}\mathbb{E}_{\mathbf{X}\sim P}\big{[}h(\mathbf{X}-\mathbf{z})^{4}\big{]}\big{)}^{1/2}\big{(}\mathbb{E}_{\mathbf{X}\sim P}\big{[}\|\mathbf{s}_{p}(\mathbf{X})\|_{2}^{4}\big{]}\big{)}^{1/2}$
@@
To show (ii), we note that a translation-invariant kernel corresponds to choosing a constant weighting function $w(\mathbf{x})\equiv 1$ in the tilted kernel of Lemma 2, so that we can leverage the derivations in the proof of Lemma 2. Setting $w(\mathbf{x})\equiv 1$ in (28) to simplify $u_{p}(\mathbf{x},\mathbf{x})$ yields
@@
Pick any $s\in[0,1)$ and any integer $m^{\prime}\in\big{[}\frac{1}{2}n^{1-s},n\big{)}$. Define $\epsilon_{m,n}\coloneqq m^{\prime}/n$ and $m=n-m^{\prime}$. Let $C,C_{P,1},C_{P,2}$ be the constants defined in Lemma 8, and define the events
@@
and the dependence on $m^{\prime}$ in these quantities is through $m=n-m^{\prime}$. It follows from Lemma 7 and Lemma 8 applied to $\delta/2$ that the event $\mathcal{A}\coloneqq\mathcal{A}_{1}\cap\mathcal{A}_{2}$ occurs with probability $\Pr(\mathcal{A})\geq 1-\delta/2-\delta/2=1-\delta$. These events allow us to bound $D^{2}(\mathbb{X}_{n,\mathbf{z}})$ and $q_{B,1-\alpha}^{2}(\mathbb{X}_{n,\mathbf{z}})$, which then allows us to show that $D^{2}(\mathbb{X}_{n,\mathbf{z}})>q_{B,1-\alpha}^{2}(\mathbb{X}_{n,\mathbf{z}})$ for sufficiently large $\mathbf{z}$. On $\mathcal{A}$, we have
@@
$\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n,\mathbf{z}})>\hat{q}_{1-\alpha}^{B}(\mathbb{X}_{n,\mathbf{z}})\big{)}\;$ $\geq\;\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n,\mathbf{z}})>\hat{q}_{1-\alpha}^{B}(\mathbb{X}_{n,\mathbf{z}})\;|\;\mathcal{A}\big{)}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}(\mathcal{A})$
@@
[…]
@@
[…]
@@
[…]
@@
A.1.4 Proof of Theorem 1
It suffices to show the claim for any sequence of the form $\epsilon_{n}=n^{-s}$, where $s\in[0,1)$ is arbitrary. Let the random variable $M^{\prime}$ be defined as at the start of Section A.1. For any random sample $\mathbb{X}_{n}$, define the event $\mathcal{A}(\mathbb{X}_{n})=\{D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\}$. For $n$ sufficiently large so that $n^{(1-s)/2}>2\sqrt{2}C\log(1/\alpha)$, we can apply Proposition 9 to conclude that, for all $n$, there exists $\mathbf{z}_{n}\in\mathbb{R}^{d}$ such that the following holds for all $m^{\prime}\in\big{[}\frac{1}{2}n^{1-s},n\big{)}=\big{[}\frac{1}{2}\epsilon_{n}n,n\big{)}$,
@@ A.1.4 Proof of Theorem 1
[…]
@@
where in the last line we have defined $f(n)=\exp\left(-\epsilon_{n}n/8\right)-\exp\left(-3\epsilon_{n}n/7\right)$.Define the index set $I_{n}$ to be the set of integers in the interval $[\epsilon_{n}n/2,3\epsilon_{n}n/3]$. Then
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\;$ $\geq\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathcal{A}(\mathbb{X}_{n})\cap\mathcal{B}\big{)}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\;\sum_{m^{\prime}\in I_{n}}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathcal{A}(\mathbb{X}_{n})\;|\;M^{\prime}=m^{\prime}\big{)}\Pr(M^{\prime}=m^{\prime})$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\;\min_{m^{\prime}\in I_{n}}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathcal{A}(\mathbb{X}_{n})\;|\;M^{\prime}=m^{\prime}\big{)}\Pr_{M^{\prime}}(\mathcal{B})$
[…]
where the first inequality follows from the law of total probability, $(a)$ and $(b)$ hold by the partition $\mathcal{B}=\cup_{m^{\prime}\in I_{n}}\{M^{\prime}=m^{\prime}\}$ and again the law of total probability, $(c)$ follows from (24) and (26), and $(d)$ holds since the assumptions $\epsilon_{n}=n^{-s}$ and $s\in[0,1)$ imply $f(n)=o\big{(}n^{-(1-s)}\big{)}$. On the other hand, the assumed moment conditions imply $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ as argued in the paragraph before (11) and $\mathbb{E}_{\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2}\sim P}[u_{p}(\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2})^{2}]\leq\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\mathbf{X}^{\ast})^{2}]<\infty$ as argued in Section A.1.3, so we can apply Arcones and Gine, (1992, Theorem 3.5) with $r=2$ to yield the asymptotic validity of the KSD test when $Q=P$, namely

### v4

@@ A.1.2 Proof of Lemma˜8
Proof of Lemma˜10  Adamczak, (2015, Remark 2.2) shows that multinomial random vectors satisfy the convex concentration inequality (Adamczak,, 2015, Definition 2.2). Therefore, we can apply Adamczak, (2015, Theorem 2.5) to conclude that there exists constants $C_{1},C_{2}$ such that, for any $n\times n$ matrix $A$ and any $t>0$,
@@
$\delta\;\geq\;2\exp\left(-\frac{1}{C_{1}}\min\left(\frac{n^{4}t^{2}}{C_{2}\|A\|_{\mathrm{F}}^{2}},\;\frac{n^{2}t}{\|A\|_{\mathrm{F}}}\right)\right)$
@@
We are now ready to prove Lemma˜8.
Proof of Lemma˜8  Denote by $F_{\infty}$ the cumulative distribution function for the conditional distribution of the bootstrap sample $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$ (defined in (3)) given $\mathbb{X}_{n,\mathbf{z}}$, and denote its $(1-\alpha)$-quantile as
$q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n,\mathbf{z}})\;\coloneqq\;\inf\{t\in\mathbb{R}:\;1-\alpha\leq F_{\infty}(t)\}\;.$
To bound $q^{2}_{\infty,1-\alpha/2}(\mathbb{X}_{n,\mathbf{z}})$, we use a similar argument as in Schrab et al., (2023, Appendix E4) by bounding the exceedance probability of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$. Notably, their proof cannot be applied directly here, because they use a wild bootstrap (Leucht and Neumann,, 2013) that corresponds to replacing $\mathbf{W}_{i}-1$ with independent Rademacher weights, whereas in our case $W_{i}$ are correlated and multinomially distributed. To proceed, we first note that we can write $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$ as a quadratic form as $D_{\mathbf{W}}^{2}(\mathbb{X}_{n,\mathbf{z}})=n^{-2}\mathbf{W}^{\top}A\mathbf{W}$, where $A=(a_{ij})_{ij}$ with $a_{ij}\coloneqq u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$. We can then use the Hanson-Wright inequality stated in Lemma˜10 applied with $\delta=\alpha/2$ to show that there exist positive constant $C$ such that, for any almost sure realizations $\mathbb{X}_{n,\mathbf{z}}$,
@@
where the last quality holds since by assumption $\mathbf{X}_{i}=\mathbf{z}$ for $i>m$. Moreover, since $a_{ii}=u_{p}(\mathbf{X}_{i},\mathbf{X}_{i})\geq 0$ for any $i$, we can apply Markov inequality to bound the first term on the RHS. This gives that the following holds with probability at least $1-\delta/4$,
@@
where the second step holds since by assumption $\mathbf{X}_{i}=\mathbf{X}_{i}^{\ast}\sim P$ are i.i.d. for $i=1,\ldots,m$, and where we have substituted $\xi_{P}^{2}=\mathbb{E}_{\mathbf{X}_{1}\sim P}[u_{p}(\mathbf{X}_{1},\mathbf{X}_{1})^{2}]$. Combining with (15) implies that, with probability at least $1-\delta/4$,
@@
$\Sigma_{P}^{2}\;=\;\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]\;\leq\;\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X})u_{p}(\mathbf{X}^{\prime},\mathbf{X}^{\prime})]\;$ $\stackrel{{\scriptstyle(a)}}{{=}}\;\big{(}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{X})]\big{)}^{2}$
@@
where $(a)$ holds since $\mathbf{X},\mathbf{X}^{\prime}$ are i.i.d. and $(b)$ follows from Jensen’s inequality. By applying Markov’s inequality again, with probability at least $1-\delta/4$, where the probability is taken over the randomness of $\mathbb{X}_{m}^{\ast}$, we have
@@
$q_{\infty,1-\alpha}^{2}(\mathbb{X}_{n,\mathbf{z}})$
$\leq\;\frac{4}{n\delta}\xi_{P}+\frac{n-m}{n^{2}}u_{p}(\mathbf{z},\mathbf{z})$
@@
A.1.3 Proof of Proposition˜9
We first show that the moment conditions (i)-(ii) in Lemma˜7 and Lemma˜8 are met, and apply these lemmas to conclude the proof.
@@ A.1.3 Proof of Proposition˜9
$u_{p}(\mathbf{x},\mathbf{z})^{2}$
$=\;\big{[}\mathbf{s}_{p}(\mathbf{x})^{\top}\mathbf{s}_{p}(\mathbf{z})h(\mathbf{x}-\mathbf{z})-\mathbf{s}_{p}(\mathbf{x})^{\top}\nabla h(\mathbf{x}-\mathbf{z})+\mathbf{s}_{p}(\mathbf{z})^{\top}\nabla h(\mathbf{x}-\mathbf{z})+\nabla^{\top}\nabla h(\mathbf{x}-\mathbf{z})\big{]}^{2}$
@@
$\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}\big{[}\|\mathbf{s}_{p}(\mathbf{X})\|_{2}^{2}\|\mathbf{s}_{p}(\mathbf{z})\|_{2}^{2}h(\mathbf{X}-\mathbf{z})^{2}\big{]}$
$\leq\;\sup_{\mathbf{z}\in\mathbb{R}^{d}}\|\mathbf{s}_{p}(\mathbf{z})\|_{2}^{2}\big{(}\mathbb{E}_{\mathbf{X}\sim P}\big{[}h(\mathbf{X}-\mathbf{z})^{4}\big{]}\big{)}^{1/2}\big{(}\mathbb{E}_{\mathbf{X}\sim P}\big{[}\|\mathbf{s}_{p}(\mathbf{X})\|_{2}^{4}\big{]}\big{)}^{1/2}$
@@
To show (ii), we note that a translation-invariant kernel corresponds to choosing a constant weighting function $w(\mathbf{x})\equiv 1$ in the tilted kernel of Lemma˜2, so that we can leverage the derivations in the proof of Lemma˜2. Setting $w(\mathbf{x})\equiv 1$ in (28) to simplify $u_{p}(\mathbf{x},\mathbf{x})$ yields
@@
Pick any $s\in[0,1)$ and any integer $m^{\prime}\in\big{[}\frac{1}{2}n^{1-s},n\big{)}$. Define $\epsilon_{m,n}\coloneqq m^{\prime}/n$ and $m=n-m^{\prime}$. Let $C,C_{P,1},C_{P,2}$ be the constants defined in Lemma˜8, and define the events
@@
and the dependence on $m^{\prime}$ in these quantities is through $m=n-m^{\prime}$. It follows from Lemma˜7 and Lemma˜8 applied to $\delta/2$ that the event $\mathcal{A}\coloneqq\mathcal{A}_{1}\cap\mathcal{A}_{2}$ occurs with probability $\Pr(\mathcal{A})\geq 1-\delta/2-\delta/2=1-\delta$. These events allow us to bound $D^{2}(\mathbb{X}_{n,\mathbf{z}})$ and $q_{B,1-\alpha}^{2}(\mathbb{X}_{n,\mathbf{z}})$, which then allows us to show that $D^{2}(\mathbb{X}_{n,\mathbf{z}})>q_{B,1-\alpha}^{2}(\mathbb{X}_{n,\mathbf{z}})$ for sufficiently large $\mathbf{z}$. On $\mathcal{A}$, we have
@@
$\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n,\mathbf{z}})>\hat{q}_{1-\alpha}^{B}(\mathbb{X}_{n,\mathbf{z}})\big{)}$
$\geq\;\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n,\mathbf{z}})>\hat{q}_{1-\alpha}^{B}(\mathbb{X}_{n,\mathbf{z}})\;|\;\mathcal{A}\big{)}\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}(\mathcal{A})$
@@
$\geq\;(1-\delta)\mathbbm{1}\bigg{\{}\epsilon_{m,n}u_{p}(\mathbf{z},\mathbf{z})\bigg{(}\epsilon_{m,n}-\frac{C^{\prime}}{n\sqrt{\delta}}\bigg{)}-T(m^{\prime})\;>\;0\bigg{\}}\;,$ (21)
@@
$n^{\frac{(1-s)}{2}}\;>\;4C^{\prime}\;,\qquad\qquad\|\mathbf{s}_{p}(\mathbf{z})\|_{2}^{2}\;>\;\frac{1}{h(0)}\Big{(}4T(n-1)n^{2s}+\nabla_{1}^{\top}\nabla_{2}h(0)\Big{)}\;.$ (22)
@@
$\Pr\nolimits_{\mathbb{X}_{m}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n,\mathbf{z}})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n,\mathbf{z}})\big{)}\;\geq\;1-\delta\;=\;1-n^{-(1-s)}\;.$
@@
A.1.4 Proof of Theorem˜1
It suffices to show the claim for any sequence of the form $\epsilon_{n}=n^{-s}$, where $s\in[0,1)$ is arbitrary. Let the random variable $M^{\prime}$ be defined as at the start of Section˜A.1. For any random sample $\mathbb{X}_{n}$, define the event $\mathcal{A}(\mathbb{X}_{n})=\{D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\}$. For $n$ sufficiently large so that $n^{(1-s)/2}>2\sqrt{2}C\log(1/\alpha)$, we can apply Proposition˜9 to conclude that, for all $n$, there exists $\mathbf{z}_{n}\in\mathbb{R}^{d}$ such that the following holds for all $m^{\prime}\in\big{[}\frac{1}{2}n^{1-s},n\big{)}=\big{[}\frac{1}{2}\epsilon_{n}n,n\big{)}$,
@@ A.1.4 Proof of Theorem˜1
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n}^{\ast})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n}^{\ast})\big{)}\;\to\;1-\alpha\;,$ (25)
@@
where in the last line we have defined $f(n)=\exp\left(-\epsilon_{n}n/8\right)-\exp\left(-3\epsilon_{n}n/7\right)$. Define the index set $I_{n}$ to be the set of integers in the interval $\epsilon_{n}n/2,3\epsilon_{n}n/3]$. Then
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}D^{2}(\mathbb{X}_{n})>q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}$
$\geq\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathcal{A}(\mathbb{X}_{n})\cap\mathcal{B}\big{)}$
$\stackrel{{\scriptstyle(a)}}{{=}}\;\sum_{m^{\prime}\in I_{n}}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathcal{A}(\mathbb{X}_{n})\;|\;M^{\prime}=m^{\prime}\big{)}\Pr(M^{\prime}=m^{\prime})$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\;\min_{m^{\prime}\in I_{n}}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathcal{A}(\mathbb{X}_{n})\;|\;M^{\prime}=m^{\prime}\big{)}\Pr_{M^{\prime}}(\mathcal{B})$
$\stackrel{{\scriptstyle(c)}}{{\geq}}\;\big{(}1-n^{-(1-s)}\big{)}\big{(}1-f(n)\big{)}\;\stackrel{{\scriptstyle(d)}}{{=}}\;1-\mathcal{O}\big{(}n^{-(1-s)}\big{)}\;,$ (27)
where the first inequality follows from the law of total probability, $(a)$ and $(b)$ hold by the partition $\mathcal{B}=\cup_{m^{\prime}\in I_{n}}\{M^{\prime}=m^{\prime}\}$ and again the law of total probability, $(c)$ follows from (24) and (26), and $(d)$ holds since the assumptions $\epsilon_{n}=n^{-s}$ and $s\in[0,1)$ imply $f(n)=o\big{(}n^{-(1-s)}\big{)}$. On the other hand, the assumed moment conditions imply $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ as argued in the paragraph before (11) and $\mathbb{E}_{\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2}\sim P}[u_{p}(\mathbf{X}^{\ast}_{1},\mathbf{X}^{\ast}_{2})^{2}]\leq\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\mathbf{X}^{\ast})^{2}]<\infty$ as argued in Section˜A.1.3, so we can apply Arcones and Gine, (1992, Theorem 3.5) with $r=2$ to yield the asymptotic validity of the KSD test when $Q=P$, namely